Pina Colada Corp. just began business and made the following four inventory purchases in June:

June   1 171 units $1026
June   10 228 units 1596
June   15 228 units 1824
June   28 171 units 1539
$5985


A physical count of merchandise inventory on June 30 reveals that there are 228 units on hand. Using the average cost method, the amount allocated to the ending inventory on June 30 is

$1710.

$1824.

$1597.

$1995.
Answer #1

Total units = 171+228+228+171 = 798 units

Average cost per unit = Total cost/Total units

= 5985/798 = 7.50 per unit

Cost of ending inventory = 228 units * 7.50 per unit

= 1710

Option A is the answer